25 A U.S.-based company, American Construction, Inc., arranged a 2-year, $1,000,000 loan to fund a project in Mexico. The loan is denominated in Mexican pesos, carries a 10.0% nominal rate, and requires equal semiannual payments. The exchange rate at the time of the loan was 5.75 pesos per dollar, but it dropped to 5.35 pesos per dollar before the first payment came due. The loan was not hedged in the foreign exchange market. Thus, Stewart must convert U.S. funds to Mexican pesos to make its payments. If the exchange rate remains at 5.35 pesos per dollar through the end of the loan period, what nominal interest rate will Stewart end up paying on the loan? Enter your answer rounded to two decimal places. Do not enter % in the answer box. For example, if your answer is 0.12345 or 12.345% then enter as 12.35 in the answer box.
Answer #1

Loan Amount= 5.75*1000000=5750000 pesos

Semi annual interest rate= 10%/2=5%

Semi annual Interest Amount= 5750000*5%=287500 pesos or 287500/5.35=$53738.32

So, semi annual interest %= 53738.32/1000000=5.374%

So, Nominal Interest Rate= 5.374%*(12/6)=10.7477% or 10.75% (Approx)

FYI, Effective Annual Interest Rate would be =(1+5.374%)^2-1=11.03%

Hence, nominal interest rate is 10.75